OpenStudy (anonymous):

Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

y-4 = 3/4 * (x+8) => y-4 = 3x/4 + 6 => y = 3x/4 + 6 + 4 => y = 3x/4 + 10

OpenStudy (anonymous):

Thank yu so much